Adaptive Parallelism Using Linda

Peter G. Clayton

Department of Computer Science Rhodes University, South Africa

Abstract

Adaptive parallelism refers to the parallel execution of a computation on a dynamically changing set of processors. The Linda coordination language provides a framework in which is it possible to implement adaptive systems on heterogeneous clusters of computers. This seminar will describe recent experiences at Rhodes University in a project to harness the ”idle cycles” of networked computers using this approach, without negatively affecting the tasks which constitute the normal workload of the computers.

Refreshments and cookies will be provided